Where to Find André DeHon

Physical

Locale  Note
East Campus-Goodale 312 Dorm Room
NE43-768A  Tech Square Office
NE43-770  Transit Lab
NE43-7th floor general  Tech Square
W20-557  SIPB Office

Virtual

Machine  Note
charon.mit.edu  SIPB time sharing machine
meyer.ai.mit.edu HP Unix box in my Tech Square Office
laplace.ai.mit.edu LISPM (3640) in my Tech Square Office
pigpen.ai.mit.edu LISPM (3640) in HQM’s Office

Email

Contex Address
Athena amdehon@athena.mit.edu
AI Lab andre@ai.mit.edu

Phone

Locale  Number Notes
Dorm Room  x5-6387 Don’t call between 5:30pm and 7pm
Tech Square Office x3-2657
SIPB Office  x3-7788 Can almost always leave a message here